OpenStudy (hippoharry):

Which of the binomials below is a factor of this trinomial? x2 - 13x + 36 A. x - 6 B. x + 6 C. x + 9 D. x - 9

OpenStudy (welshfella):

to find the factors you need to get 2 numbers a and b where ab = 36 and a + b = -13

OpenStudy (welshfella):

check out the factors of 36 6 and 6 -6 and -6 -9 and -4 9 and 4 would one of these fit?

OpenStudy (welshfella):

- they need to equal -13 when added

OpenStudy (welshfella):

the the factor will be (x + a) where a is one of the numbers

OpenStudy (welshfella):

a simple example would be x^2 + 4x + 4 we see that 2*2 = 4 and 2 + 2 = 4, so the factors are (x + 2)(x + 2)

OpenStudy (hippoharry):

but that doesnt make sense for the answers it gave me to choose @welshfella

OpenStudy (welshfella):

yes it does make sense the reqiured numbers are -4 and -9 because they multiply to + 36 and add to -13 so the factors are (x - 9)(x - 4) look thru the list and the one you want is (x - 9)

OpenStudy (welshfella):

follow?

OpenStudy (welshfella):

you put all the effort in to explain and they just disappear!!